oThe amount of calcium and iron (in milligrams) found in one 10 oz package of broccoli
If there 13% of calcium in one 10oz package of broccoli and 12% of iron how do I convert percentage to milligrams?

Answer by Alan3354(69443)

If there 13% of calcium in one 10oz package of broccoli and 12% of iron how do I convert percentage to milligrams?
---------------
1 ounce =~ 28.35 grams = 28,350 milligrams
10 ounces = 283,500 milligrams
------------------
13% of 283500 = 0.13*283500 = 36,855 milligrams of Ca
12% = 0.12*283500 = 34,020 milligrams of Fe
-------------------
Package labels usually show some %age of MDR (Minimum Daily Requirements) of things, not the actual contents. I doubt that broccoli is 13% calcium.